GreenHab Report – February 5th

GreenHab Officer: Mehnaz Jabeen
Environmental control: heater on, fan off, door closed,
Average temperatures:
5:42 : 73.2F, 25%
Hours of supplemental light: None
Daily water usage for crops: 8 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank (200 gallon capacity): 192 gallons
Time(s) of watering for crops: 5:44
Changes to crops: None
Narrative: Today was the first official sol as greenhab officer. Right after the training, which got over around 17:30, I went to the hab with Clare, our crew journalist who helped me water the plants and briefed about the crops and activities from the previous crew as she took part in the previous expedition. As soon as we got inside the heater turned on. All the plants looked good except for the cucumbers, the leaves were drying out. Therefore I’m aiming to water them a bit more ( 1 gallon) than the previous crew did daily. I also found one of the big blue containers with no growing seed so I would like to plant some new crops. Many tomatoes were ripening and were red, almost on the last stage so after one or two days they shall be ready for harvest.

Harvest: None
Support/supplies needed: Seeds : zucchini , cauliflower, cabbage

Astronomy Report – February 5th



Name: Aditya Krishna Karigiri Madhusudhan
Crew: 292
Date: 05/02/2024

Robotic Telescope Requested: MDRS-WF
Objects to be Imaged this Evening: NGC 281 (Pacman Nebula). Observation submitted in Skynet on 05/02/2024 at 6:33 pm
Images submitted with this report: None

Problems Encountered: None

Solar Features Observed: None
Images submitted with this report: Please find the images of the astronomy laptop box attached to this email.
Problems Encountered: I went to the Solar Musk Observatory at 6 pm on 05/02/2024 to take images of the telescope and the astronomy box. I used the code provided to open the lock but the lock didn’t open. Please find a video showing the issue. I tried wiggling the lock several times to see if it was stuck somewhere but couldn’t succeed. I will give it a try again tomorrow morning.

HSO Beginning Of Mission Checklist – February 5th

Submitted by: Daniel Loy

Crew: 292

Date: 05/02/2024

Part 1: Using the attached Safety Equipment Inventory, locate, test and confirm operation of all safety equipment. List any equipment not found and/or missing:

Channel 10 radio in the science dome suspected to have dead battery, as turns on and functions when plugged in and charging, but when taken out of the charger it is non-functional.

Fire extinguishers: 3 are overdue maintenance inspections, HAB Upper and Science Dome last checked 19/01/2023, HAB Lower last checked 01/01/2023. RAM and GreenHab in date and last inspected 05/01/2024.

Part 2: Locate and confirm the emergency escape routes in the Hab are functional and clear:

  1. Stairs

  2. Emergency window

  3. Commander’s window

All escape routes checked and are functional and clear for use.

Part Three:

Inventory First Aid kit and note what needs to be refilled:

HAB Lower First aid kits

IFAK Trauma Kit Inventory list did not exist in the kit so made the below list of contents:

1 pair large gloves

5mx2.5cm adhesive tape

1 tourniquet

4.5mx5cm elastic bandage

1 mini sharpie

40x60cm dressing

1 triangular bandage

3 inch splint

1 compressed z-folded gauze

40 alcohol prep pads

1 emergency thermal blanket

1 pair scissors

1 emergency compression bandage

Also one plastic bag containing: 10 adhesive bandages 7.6×1.9cm, 4 adhesive bandages 7.6×4.5cm, 2 5x5cm sterile gauze swab and 2 antiseptic towelette.

Large first aid cabinet had no list so the following inventory was created:

3 first aid books

4 inch sports wrap

Open box of cotton swabs over ¾ full

2×4.1 inch stretch bandage roll

Open box of masks, over 15 remaining

1 x pen

Infrared thermometer

Blood pressure monitor

Finger pulse oximeter

2x isopropyl alcohol, 473 and 946ml, both open and over ¾ full

473ml hydrogen peroxide, over ¾ full

10+ 2 pack ibuprofen

This also contained a smaller first aid box that contained:

1 first aid guide

2 x thermal blanket

1 x plastic tweezers

3 x triangular bandage and pin

1 Dramamine motion sickness vial

1 Dramamine motion sickness pill

2 x small gauze wrap

4x7cm elastic bandage

Partial used durapore wrap

Partial used zinc oxide tape

Science dome first aid kit inventory as crew 291 left it with list, this includes:

1 pack gloves

1 triangular bandage

3 rolls gauze bandages

1 rescue blanket

1 pack safety pins

1 4×4 inch sterile pad

1 pair of scissors

16 alcohol swabs

Also 2 x eyewash, one expired October 2021, one with expiry date December 2024.

OSHA First aid kit had discrepancies as listed below in missing health and safety supplies section

Note any safety issues:

Note any health/environmental issues:

Note any missing or recommended health and safety supplies:

  • Only 8 1×3 inch adhesive bandages, lower than minimum fill of 16.

  • No adhesive tape.

  • Only 6 packs antibiotic ointment, lower than minimum fill of 10.

  • Only 8 packs burn treatment, lower than minimum fill of 10.

  • 0 sterile pads.

  • 0 trauma pads.

Safety Equipment Inventory 2023-2024

HAB Upper deck

HAB Lower deck





CO monitor






Escape ladder




Fire blanket




Fire extinguisher






First Aid











Propane alarm



Radios (Channels 10 and 22)






Smoke alarm






Tow rope


Sol Summary – February 5th

Summary title : Mangalyatri, Crew 292 Begins Simulation After Training.
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Activity Summary: Despite delays and snowy stormy weather Mars Society Australia Crew 292 Mangalyatri (Mars Explorers) all finally arrived at the Mars Desert Research Station, joining crew member Clare Fletcher who was also a member of MSA Crew 291. This morning we trained for simulation and learnt about all the systems of the Mars hab. We spent time in the engineering bay (RAM – the repair and maintenance module), in the fruiting and plentiful Green Hab and in the extraordinary Science Dome. Later, outside, MDRS Director Sergii Lakymov taught us how to charge and drive the electric rovers and we all went for a practice drive. For our crew geologist Bharti Sharma, it was her first experience ever behind the wheel of a vehicle. In the afternoon we were trained in comms with radios and headsets and then we tried on space suits. This was great and congratulations, North California Mars Society – the suits you have made for the station are excellent, so comfortable and easy to wear. At 3.30 pm everyone sat down for a cup of tea and a quick planning meeting. We discussed tomorrow’s two training EVA’s, and our dinner roster for this evening. We examined our water use and thought about how we can conserve our resources. We scheduled a research discussion meeting for tomorrow and decided that we would begin simulation at six pm today. With two hours to go of fresh air outside, we all went for a walk up to Hab Ridge. The Henry Mountains are snow-covered and beautiful and we wandered along the top of the Dakota sandstone. We came back to the hab via spectacular North Ridge with its deco colours – creamy pinks, chocolate browns and ochres, rusts and black. The desert was quiet and still, the ground soft from the rain and we walked home on pink mud like Mars. We came inside, shut the doors and began simulation.
Look Ahead Plan: Two training EVA’s and a research project meeting to roster in field work.
Anomolies in work: Nothing to report
Weather: Light cloud cover today with mild wind, NNE 8-17 mph. Daily temperature 51-54F.
Crew Physical Status: Nominal
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, EVA Request(s)
Support requested: Nothing to request.

Operations Report – February 5th

SOL: #1 (from 6:00 pm)
Name of person filing report: Rajvi Patel
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit Rover used: Not used
Opportunity rover used: Not used
Curiosity rover used: Not used
Perseverance rover used: Not used
General notes on rovers: Training day today
Summary of Hab operations:
WATER USE: # gallons (Reference depth not available for yesterday)
Water (static tank): 462.56 gallons (Depth = 6.69 inches)
Static tank pipe heater (on or off): on
Static tank heater (On or off): on
Toilet tank emptied: yes
Summary of internet: Nominal during comms window
Summary of suits and radios: Nothing to report for today.
Summary of GreenHab operations:
WATER USE: 8 gallons
Heater: On
Supplemental light: 2200- 0200
Harvest: None
Summary of Science Dome operations: None
Dual split: Heat On
Summary of RAM operations: (Tools used, work done)
Used a measuring tape and cleaned it, to measure the water depth
Summary of any observatory issues: The crew astronomer could not open the solar observatory due to the lock on it. (Check out Crew Astronomer’s report)
Summary of health and safety issues: Check out the HSO report for details
Questions, concerns, and requests to Mission Support: None
Hab Car exit Mileage: 220571 miles

Journalist Report – February 5th

Sol 1: Welcome to Mars!
Well, Crew 292 Mangalyatri have officially landed on Mars! Many of the Crew had a challenging journey here, while I waited for them after Crew 291’s departure back to Earth, but everyone got here safe and sound in the end. We had a big dinner to celebrate our landing and an even bigger sleep to feel refreshed before our training began this morning.
Sol 1 began with coffee, tea, and delicious banana pancakes – the perfect breakfast to fuel a big sol of training and our sim beginning.
We started our training at 0900 and kept training until about 1500 (though we did stop for a much-needed tea break and lunch break). We’ve all learned a lot of valuable skills about how to do our jobs around the Hab and how to drive, use EVA suits, and use radios. We were all feeling prepared to start sim but couldn’t resist taking lots of photos and videos and going for a big walk before we could no longer take walks without EVA suits and a strong scientific justification. Crew 292 greatly enjoyed going up to Hab Ridge and looking for fossilised oyster shells, before heading back to the Hab about 40 minutes before our sim began.
We finished off some last-minute chores like watering plants and checking our supplies before it ticked over to 1800 and we were officially living on Mars.
All that’s left to complete our first sol on Mars is a delicious and nutritious dinner, getting used to our schedule of report writing, communications, cooking, and working, and having a big sleep before beginning our EVAs tomorrow. All our Mangalyatris are feeling welcome and settled and are super excited to begin our rotation on Mars!
Until next sol,
Clare Fletcher (292 Mangalyatri journalist)

Sol Summary – February 7th

Sol: Sol #3
Summary Title: A windy but productive day for Mangalyatri Crew 292.
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Activity Summary: The weather was extreme today. Even though the desert looked still, we could hear the wind howling around the hab when we woke. We correctly anticipated a delay in our fieldwork. Most of us spent the morning organizing field materials and doing research. Mid-morning Bharti and I assisted our engineer Rajvi to check the covers for the tunnels, directly addressing any issues. We made sure all our doors are well-secured and that we are keeping good radio contact in the different working zones of the hab. Mehnaz did her Green Hab chores early and then cooked aloo sandwiches (spicy potato) with Bharti. We made some yoghurt in the breadmaker, which is fantastic, and we cooked chick-peas in the slow cooker for tonight’s dinner. Late morning, yoga and we all had a late lunch. Straight after lunch, in a quick group meeting, Bharti demonstrated her idea for saving water while washing our dishes and we all agreed to change dish washing practices. Because of the high winds, as predicted, our EVA this afternoon was postponed. However, we made the best of our time inside. This afternoon everyone worked on their own projects. Aditya set up an observation in the robotic telescope and then did his analysis. Clare and I spent time in the science dome and set up a working table for drawing and painting samples, in this case, gypsum. Daniel extracted 0.3 grams of salt from a gypsum crystal to put into different media for his samples. Mehanz had a useful discussion with Rajvi and Aditya about her research and returned to the Green Hab to harvest radish, coriander and basil and check the moisture levels of the plants. Bharti read papers sent by Dr. Jon Clarke on slope analysis and basic planetary geology. Dinner tonight is chana masala cooked by me (chickpeas with tomatoes and onions) and paratha, (flat bread) freshly made by Clare. (BTW the Indian food cooked by everyone on this crew is excellent). Maybe another game of cards after Comms and fingers crossed again for good weather tomorrow. We are ready for fieldwork.
Look Ahead Plan: Today’s postponed EVA to Cowboy Corner requested for tomorrow.
Anomalies in work: None
Weather: Very windy with SSW gusts of 25mph. The forecast today was cool with high cloud cover, maximum temperature 50F, now dropping.
Crew Physical Status: Nominal
EVA: none
Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA report, Astronomy Report, EVA request.
Support Requested: None

Operations Report – February 7th

SOL: #3
Name of person filing report: Rajvi Patel
Non-nominal systems: None
Notes on non-nominal systems: None
Spirit Rover used: Not used
Opportunity rover used: EVA #3 was postponed due to the windy conditions.
Curiosity rover used: EVA #3 was postponed due to the windy conditions.
Perseverance rover used: Not used
General notes on rovers:
Summary of Hab operations: put notes here
WATER USE: 23.2 gallons
Water (static tank): 416.21 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes
Summary of Internet: Nominal
Summary of suits and radios: Not used
Summary of GreenHab operations: put notes here
WATER USE: 7 gallons
Heater: ON
Supplemental light: 22:00 – 02:00
Harvest: 117g of radish, 23g of Coriander + Basil
Summary of ScienceDome operations: Crew Biologist transferred brought equipment and consumables to the ScienceDome for their project. 140ml of 95% Ethyl Alcohol was used with 60ml of DI water to make 200ml of 70% ethanol for surface sterilization. The fume cabinet was cleaned to make two 5.5ml 1:10 dilutions of brought Flour, SP, and LB media. One of these was kept as a negative control while the other was inoculated with 0.1 grams of gypsum extracted from a larger gypsum crystal from the corner of Brahe Highway and Cow Dung Road from a crew 291 EVA. These were then kept in Oven 1, not turned on, to keep the samples at the internal temperature of 28º C.
Equipment used: Ethanol spray bottle, tweezers, weighing scale (new batteries put in), fume cabinet, cleaning tissues, Oven 1 (not turned on), 100-1000ul pipette, and own pipette tips.
Crew Artist and Journalist used the Science Dome for drawing and painting rock samples for their projects.
Dual split: Heat ON
Summary of RAM operations: The crew Engineer used it to get the tools wire cutter and heavy-duty zip locks to secure locations on the tarp using zip ties. Tools used were wire cutter and heavy-duty zip ties. Used measuring tape to measure the water level. Borrowed another measuring tape for the Crew Biologist’s project in Science Dome.
Now secured
Summary of any observatory issues: None
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: put notes here
-Most of the locations were secured using heavy-duty zip ties. However, there are some locations (with torn holes for zip ties to go through) that are required to be secured. (Picture 1 and 2). Advice?
-The photo about securing the bottom of the overlapping tarp section mentioned in the last report is also attached below (highlighted in yellow), along with the type of secured arrangement I would like to propose (highlighted in blue). Advice?

Journalist Report – February 7th

Sol 3: Food, glorious food! Wind, horrible wind!
The crew had a slow morning today, thinking about our research and doing a few maintenance jobs around the Hab, various domes, and tunnels. We were all feeling a little anxious about the weather and what that would mean for the EVA we had planned for this afternoon. It wasn’t long before we decided that being in the tunnels during the high winds was both unpleasant and dangerous in the thin Martian atmosphere.
To take our mind off the possibility of the afternoon’s EVA getting postponed due to high winds, we cooked aloo sandwiches. And lots of them! It was a real team effort, with everyone taking part in the preparation, tasting, and actual cooking (and of course eating!) of these warm and comforting spiced mashed potato sandwiches. As we all ate, we discussed reality TV, what “going on a spiritual journey” means, and more serious topics of politics and religion.
By 1330 we decided to radio back to Mission Support and say that we would postpone our EVA until tomorrow, and Mission Support had recommended that we do this if the wind didn’t die down. It was just too windy to be out and about on rovers and in space suits. We’d all gotten sick of hearing the wind howling around the Hab, but not being able to see any sign of said wind when we looked outside.
So, we all settled in for an afternoon of reading and researching. Three Mangalyatris went to the Science Dome – one preparing for his forthcoming biological experiments, and the other two to look at, draw, and paint gypsum and chert as part of Annalea’s research. Our GreenHab Officer enjoyed the warmer temperatures in the GreenHab while taking care of our plants and coming to grips with her research about how different conditions affect plant germination and growth. She also harvested some radish, basil, and coriander – what a treat for dinner tonight!
Once all the day’s tasks were over, our evening became one of cooking, report writing, and continued research. All our cooking is keeping spirits up while hoping for the weather to calm down. Plus, it fuels all our important research! We’re writing and researching as our yoghurt sets, our chickpeas soak, and our paratha dough rests. After a full day of inside work, the Mangalyatri crew is looking forward to getting out on an EVA. But perhaps we are even more looking forward to a post-comms window board game or cards.
Hoping for better weather and the same quality of food,
Clare Fletcher (292 Mangalyatri Journalist)

Copyright © The Mars Society. All rights reserved. | Main Site